You are scuba diving underwater (n = 1.33) near an
Answer #1

Part a) according to the snell's law we know that if light is passing from one medium of refractive μ1 to another medium of refractive index μ2 then according to the law

μ1 sin i= μ2 sin r

it is given in the question that light is passing from water to diamond

therefore our μ1water= 1.33

and μ2diamond= 2.42

and angle of incidence is given as i=25o

μ1 sin i= μ2 sin r

1,33 sin25o= 2.42 sin r

sin r=1.33 sin25o/2.42

sin r=(1.33 *0.4226182)/2.42

sinr=0.232265

r= sin-10.232265

r=13.430 o

This is the angle at which ray refract inside diamond

Part B) in order to make the light reflect back to water itself we need to have total internal reflection

therfore angle of refratio would be 900 which is nothing but the critical angle

according to snells law

μ1 sin i= μ2 sin r

μ1 =μwater= 1.33

μ2 =μair= 1

sin r=sin 900=1

therfore 1.33 sini =1 *1

sin i=1/1.33

sini=0.751879

i= sin-10.751879

i=48.750

This is the minimum angle at which ray reflect back to water

Part c) speed of light in air=3 * 108 m/s

speed of light in any medium is given by =speed of light in air/refractive index of the medium

speed of laser beam in diamond=3 * 108/2.42

speed of laser beam in diamond=1.23966 *108 m/s

speed of laser beam in water=3 * 108/1.33

speed of laser beam in water=2.2556 *108 m/s

GOOD LUCK!!

angle at which light will reflect back